OpenStudy (anonymous):

A bag of fertilizer covers 2500 square feet of lawn. Find how many bags of fertilizer should be purchased to cover a rectangular lawn 410 feet by 65 feet. a. 11 bags c. 1066 bags b. 10 bags d. 107 bags

OpenStudy (anonymous):

@texaschic101

OpenStudy (nirmalnema):

i think the option b is 10.66 bags and it is the correct ans..

OpenStudy (nirmalnema):

total area = 410*65=26650sqft 1 bag covers 2500 sqft so no of bags =26550/2500 =10.66 bags

OpenStudy (texaschic101):

I agree

OpenStudy (anonymous):

the correct answer is A